OpenStudy (laddiusmaximus):

derivative of (1+sectheta) is it just sectheta tantheta?

OpenStudy (dumbcow):

hmm isn't it? \[d/d \theta \sec \theta = \sec \theta \tan \theta\]

OpenStudy (calculusfunctions):

Sorry! For a secand there I thought you were talking about hyperbolic functions.

OpenStudy (calculusfunctions):

Yes you're right!

OpenStudy (laddiusmaximus):

thanks. I tend to look at 1+sec or any other trig function as one whole function and not separate parts. 1 is of course zero